# Positive Semidefinite Matrix Supermartingales

### v4

@@ 1 Introduction
for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
Here, crucially, the order $\preceq$ is the Loewner order, where $\mathbf{A}\preceq\mathbf{B}$ if and only if $\mathbf{B}-\mathbf{A}$ is a positive semidefinite matrix. This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “$=$”.
@@
for $\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
@@
    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3 and 3.4) that generalize the corresponding scalar results.
@@
    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.10), symmetric (Example 3.9), or heavy-tailed (Examples 3.8 and 3.12) martingale-dependent matrices.
@@
    Matrix e-values and e-processes (Definitions 3.13 and 3.14), generalizing the corresponding scalar concepts that have recently been extensively studied.
@@
The rest of the paper is organized as follows. We review in Section 2 the technical background of matrix analysis, sequential statistics with scalar processes in the abovementioned classes, as well as matrix concentration inequalities in the literature. We study forward PSD matrix supermartingales, their Ville’s inequality and applications in Section 3. We then study backward PSD matrix submartingales, their Ville’s inequality and applications in Section 4. In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Appendix A. An interesting empirical study is presented in Section 6, including a problem where our matrix method yields higher out-of-the-box power compared to the previous scalar method. Omitted proofs are found in Appendix B; some further references and independent theoretical results in Appendix C.
@@ 2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^1] We only briefly discuss the case with non-symmetric and non-square matrices in Section C.3. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. We consider random matrices taking values in $\mathcal{S}_{d}$. It is easy to see that if $\mathbf{X}$ is an integrable random matrix taking values in $\mathcal{S}_{d}^{I}$ where $I$ is an interval, any conditional expectation $\mathbb{E}(\mathbf{X}|\mathcal{A})$ also takes values in $\mathcal{S}_{d}^{I}$. The Loewner partial order $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^2] We use $\operatorname{abs}(\,\cdot\,)$ instead of ||$\cdot$|| ⋅ | to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$.
[…]
@@ 2.2 Scalar Nonnegative Supermartingales and Statistical Applications
As Markov’s inequality lays the foundation for a large body of concentration inequalities for a fixed number of random variables leading up to non-asymptotic statistical methods (e.g. confidence intervals) valid at fixed finite sample sizes, Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities and accordingly sequential statistical methods which are valid at data-dependent sample sizes. For example, a $(1-\alpha)$-confidence sequence, which is derived almost exclusively via (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b], is a sequence of confidence sets $\{C_{n}\}_{n\geqslant 1}$ for the parameter $\theta$ such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$ that can possibly depend on the data. More generally, the stopped version of (VI) is used in sequential hypothesis testing, as it implies that to construct a conservative sequential test of $\alpha$ level it suffices to construct a process $\{Y_{n}\}$ with $Y_{0}=1$ that is a nonnegative supermartingale under the null hypothesis, and reject it when $Y_{n}$ first surpasses $1/\alpha$. Various examples of sequential tests constructed via nonnegative supermartingales can be found in the recent review article by Ramdas et al. [2023, Sections 4 and 5]. We return to this topic of test processes with a more involved discussion in Section 3.4.
@@ 2.3.2 Ahlswede-Winter Bounds
Our time-uniform, martingale-based main results are often obtained via (MMI). In particular, we show that this Loewner-ordered approach pioneered by Ahlswede and Winter [2002] does lead to sharp matrix concentration inequalities and tests, fixed-time and time-uniform alike, comparable to and sometimes better than the results cited in Section 2.3.1. We provide a self-contained proof of (MMI) in Appendix A that also comes with the extension of randomization.
@@ 3.1 Convergence, Optional Stopping, and Maximal Inequalities
Recall our definition of a matrix-valued supermartingale from the beginning of the paper that an adapted $\mathcal{S}_{d}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under $+$). See also a “dominated integrability” lemma we prove in Section B.1.
Matrix supermartingales are related to scalar supermartingales in the following ways. First, the monotonicity of trace implies that if $\{\mathbf{Y}_{n}\}$ is an $\mathcal{S}_{d}$-valued supermartingale, its trace process $\{\mathrm{tr}(\mathbf{Y}_{n})\}$ is a real-valued supermartingale. Second, we have the following lemma.
@@
An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale.
The very straightforward proof of Lemma 3.1 can be found in Section B.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
@@
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\tau$ be a stopping time. Then,
@@
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale.
@@
    If $\tau\leqslant N$ almost surely for some $N$, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
  3. 3.
    If $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$.
@@
1. For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on $\{\mathcal{F}_{n}\}$. 2. For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{\tau})\mathbf{v}=\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v})\leqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{0}\mathbf{v})=\mathbf{v}^{\mathsf{T}}(\mathbb{E}\mathbf{Y}_{0})\mathbf{v}$. Therefore $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. 3. Applying Theorem 3.2 to the $\mathcal{S}_{d}^{+}$-valued stopped supermartingale $\{\mathbf{Y}_{\tau\wedge n}\}$, we have $\mathbb{E}\mathbf{Y}_{\tau\wedge\infty}\preceq\mathbb{E}\mathbf{Y}_{\tau\wedge 0}$, i.e., $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. ∎
@@
for any stopping time $\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
@@
We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotinicity of the trace,
@@
As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section C.2. It is not necessarily relevant to our upcoming discussion in statistical contexts but might be of independent interest.

### v6

@@ 1 Introduction
$\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}.$ (1)
Here, crucially, the order $\preceq$ is the Loewner order, where $\mathbf{A}\preceq\mathbf{B}$ if and only if $\mathbf{B}-\mathbf{A}$ is a positive semidefinite matrix. This introduces the rich spectral theory of matrices to the familiar $d=1$ scalar special case as well as the matrix martingale special case where the $\preceq$” in (1) takes the entry-wise “==”.
@@
$\text{for all }n\geqslant 1,\quad\mathbb{E}(\mathbf{Y}_{n-1}|\mathcal{G}_{n})\succeq\mathbf{Y}_{n}.$ (2)
@@
    The convergence and optional stopping theorems, and the Ville’s inequality for PSD matrix supermartingales (Theorems 3.2, 3.3, and 3.4) that generalize the corresponding scalar results.
@@
    Sequential testing procedure with a matrix test supermartingale (Algorithm 1) that can be applied to the meaning testing of bounded (Examples 3.7 and 3.11), symmetric (Example 3.10), or heavy-tailed (Example 3.9 and 3.13) martingale-dependent matrices.
@@
    Matrix e-values and e-processes (Definitions 3.14 and 3.15), generalizing the corresponding scalar concepts that have recently been extensively studied.
@@
The rest of the paper is organized as follows. We review in Section 2 the technical background of matrix analysis, sequential statistics with scalar processes in the abovementioned classes, as well as matrix concentration inequalities in the literature. We study forward PSD matrix supermartingales, their Ville’s inequality and applications in Section 3. We then study backward PSD matrix submartingales, their Ville’s inequality and applications in Section 4. In Section 5 we provide several perspectives comparing our inherent matrix approach to some existing scalar-born or scalarized results, and discuss a randomized improvement that is further developed in Section 7. An interesting empirical study is presented in Section 6, including a problem where our matrix method yields higher out-of-the-box power compared to the previous scalar method. Omitted proofs are found in Section 8; some further references and independent theoretical results in Section 9.
@@ 2.1 Matrix Analysis
Let $\mathcal{S}_{d}$ denote the set of all $d\times d$ real-valued symmetric matrices, which is the only class of matrices considered in this paper. [^2] We only briefly discuss the case with non-symmetric and non-square matrices in Section 9.3. These matrices are denoted by bold upper-case letters $\mathbf{A},\mathbf{B}$, etc; whereas $d$ dimensional (column) vectors by bold lower-case letters $\mathbf{u},\mathbf{v}$, etc. For $I\subseteq\mathbb{R}$, we denote by $\mathcal{S}_{d}^{I}$ the set of all real symmetric matrices whose eigenvalues are all in $I$. $\mathcal{S}_{d}^{[0,\infty)}$, the set of positive semidefinite and $\mathcal{S}_{d}^{(0,\infty)}$, the set of positive definite matrices are simply denoted by $\mathcal{S}_{d}^{+}$ and $\mathcal{S}_{d}^{++}$ respectively. We consider random matrices taking values in $\mathcal{S}_{d}$. It is easy to see that if $\mathbf{X}$ is an integrable random matrix taking values in $\mathcal{S}_{d}^{I}$ where $I$ is an interval, any conditional expectation $\mathbb{E}(\mathbf{X}|\mathcal{A})$ also takes values in $\mathcal{S}_{d}^{I}$. The Loewner partial order $\mathbf{A}\preceq\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive semidefinite, and $\mathbf{A}\prec\mathbf{B}$ means $\mathbf{B}-\mathbf{A}$ is positive definite.
Any function $f:I\to J$ (where $I,J\subseteq\mathbb{R}$) is identified canonically with a function $f:\mathcal{S}_{d}^{I}\to\mathcal{S}_{d}^{J}$ in the following way: for a diagonal matrix $\mathbf{\Sigma}\in\mathcal{S}_{d}^{I}$, $f(\mathbf{\Sigma})$ is obtained by applying $f$ entrywise; otherwise it is obtained by applying $f$ after orthogonal diagonalization, $f(\mathbf{U}^{\mathsf{T}}\mathbf{\Sigma}\mathbf{U})=\mathbf{U}^{\mathsf{T}}f(\mathbf{\Sigma})\mathbf{U}$. Notable matrix-to-matrix operators that recur in this paper defined this way include the matrix exponential $\exp:\mathcal{S}_{d}\to\mathcal{S}^{++}_{d}$, the matrix logarithm $\log:\mathcal{S}^{++}_{d}\to\mathcal{S}_{d}$, the matrix absolute value [^3] We use $\operatorname{abs}(\,\cdot\,)$ instead of $|\cdot|$ to distinguish it from the determinant and other norms. $\operatorname{abs}:\mathcal{S}_{d}\to\mathcal{S}^{+}_{d}$, and powers $\mathbf{X}\mapsto\mathbf{X}^{k}$ for any $k\in\mathbb{R}$.
Important functionals $\mathcal{S}_{d}\to\mathbb{R}$ we use throughout the paper include $\lambda_{\max}$, the largest eigenvalue of a matrix; $\|\cdot\|$, the spectral norm of a matrix which equals $\lambda_{\max}\circ\operatorname{abs}$; and $\mathrm{tr}$, the trace. Note that $\mathrm{tr}$ preserves the orders (viz., $\mathbf{A}\preceq\mathbf{B}$ implies $\mathrm{tr}\mathbf{A}\leqslant\mathrm{tr}\mathbf{B}$) and is lower bounded by $\lambda_{\max}=\|\cdot\|$ on $\mathcal{S}_{d}^{+}$.
@@ 2.2 Scalar Nonnegative Supermartingales and Statistical Applications
As Markov’s inequality lays the foundation for a large body of concentration inequalities for a fixed number of random variables leading up to non-asymptotic statistical methods (e.g. confidence intervals) valid at fixed finite sample sizes, Ville’s inequality (VI) has been the workhorse for many recent time-uniform concentration inequalities and accordingly sequential statistical methods which are valid at data-dependent sample sizes. These methods enable conducting hypothesis testing and estimation with uncertainty quantification without being committed to an a priori fixed sample size, allowing the experimenter to stop the experiment for any (possibly data-dependent) reason — formally, at a stopping time $\tau$ on the filtration $\mathcal{F}_{n}=\sigma(X_{1},\dots,X_{n})$ generated by the data sequence $X_{1},X_{2},\dots$.
  • A $(1-\alpha)$-confidence sequence for the parameter $\theta$ is a sequence of sets $\{C_{n}\}_{n\geqslant 1}$ (each $C_{n}$ computed from $X_{1},\dots,X_{n}$) such that $\mathbb{P}(\theta\in C_{n}\text{ for all }n)\geqslant 1-\alpha$, or equivalently $\mathbb{P}(\theta\in C_{\tau})\geqslant 1-\alpha$ for any stopping time $\tau$. They are usually derived via equation (VI) [Howard et al., 2021, Waudby-Smith and Ramdas, 2023, 2020, Wang and Ramdas, 2023a, b].
  • When testing the null hypothesis that the true data-generating distribution belongs to the set $\mathcal{H}_{0}$, a test supermartingale for $\mathcal{H}_{0}$ is a process $\{Y_{n}\}$ (each $Y_{n}$ computed from $X_{1},\dots,X_{n}$) with $Y_{0}=1$ that is a nonnegative supermartingale, i.e. $\mathbb{E}_{P}(Y_{n}|X_{1},\dots,X_{n-1})\leqslant Y_{n-1}$, under any distribution $P\in\mathcal{H}_{0}$. One rejects $\mathcal{H}_{0}$ when the observed $Y_{n}$ first surpasses $1/\alpha$. This ensures, due to (VI), that the type-I error rate of ever falsely rejecting $\mathcal{H}_{0}$ is at most $\alpha$.
These (VI)-driven sequential methods are widely applied in real-world cases where it is much more economical (see e.g. the argument made by Ham et al. [2023]) to leave the sample size unspecified and allow flexible sequential experiments. See the recent review article by Ramdas et al. [2023, Sections 4 and 5] for numerous examples. We revisit this topic with a more involved discussion on test processes in Section 3.4.
@@ 2.3.2 Ahlswede-Winter Bounds
Our time-uniform, martingale-based main results are often obtained via (MMI). In particular, we show that this Loewner-ordered approach pioneered by Ahlswede and Winter [2002] does lead to sharp matrix concentration inequalities and tests, fixed-time and time-uniform alike, comparable to and sometimes better than the results cited in Section 2.3.1. We provide a self-contained proof of (MMI) in Section 7 that also comes with the extension of randomization.
@@ 3.1 Convergence, Optional Stopping, and Maximal Inequalities
Recall our definition of a matrix-valued supermartingale from the beginning of the paper that an adapted $\mathcal{S}_{d}$-valued stochastic process $\{\mathbf{Y}_{n}\}$ is a supermartingale if $\mathbb{E}(\mathbf{Y}_{n}|\mathcal{F}_{n-1})\preceq\mathbf{Y}_{n-1}$ holds for every $n$. Before we proceed, note that the order $\preceq$ is preserved by addition, limit, and Lebesgue integration (hence including convergent infinite sums, expectations, conditional expectations, etc., essentially because $\mathcal{S}^{+}=\{\mathbf{A}:\mathbf{A}\succeq 0\}$ is closed both topologically and under ++). See also a “dominated integrability” lemma we prove in Section 8.1.
Matrix supermartingales are related to scalar supermartingales in the following ways. First, the monotonicity of trace implies that if $\{\mathbf{Y}_{n}\}$ is an $\mathcal{S}_{d}$-valued supermartingale, its trace process $\{\mathrm{tr}(\mathbf{Y}_{n})\}$ is a real-valued supermartingale. Second, we have the following lemma ($\mathbb{Q}$ being the set of rational numbers).
@@
(1) An $\mathcal{S}_{d}$-valued adapted process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if for every non-random vector $\mathbf{v}\in\mathbb{R}^{d}$ or $\mathbf{v}\in\mathbb{Q}^{d}$, the scalar-valued process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale. (2) Similar to the scalar case, an $\mathcal{S}_{d}$-valued integrable process $\{\mathbf{Y}_{n}\}$ is a supermartingale if and only if the predictable component of its Doob decomposition $\mathbf{A}_{n}=\sum_{i=1}^{n}(\mathbb{E}(\mathbf{Y}_{i}|\mathcal{F}_{i-1})-\mathbf{Y}_{i-1})$ is decreasing in the $\preceq$ order.
The proof of Lemma 3.1 can be found in Section 8.2. It leads to an analog of Doob’s martingale convergence theorem for PSD matrix supermartingales.
@@
Let $\{\mathbf{Y}_{n}\}$ be an $\mathcal{S}_{d}$-valued supermartingale, and $\sigma\leqslant\tau$ be two stopping times. Then:
@@
    $\{\mathbf{Y}_{n\wedge\tau}\}$ is also a supermartingale on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$.
@@
    If (a) $\tau<N$ for some finite $N$, or (b) $\tau<\infty$ and $\{\mathbf{Y}_{n}\}$ is $\mathcal{S}_{d}^{+}$-valued, then $\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})\preceq\mathbf{Y}_{\sigma}$. Consequently, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{\sigma}\preceq\mathbb{E}\mathbf{Y}_{0}$.
@@
The proofs of two statements are as follows.
  1. 1.
    For any $\mathbf{v}\in\mathbb{R}^{d}$, the scalar process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ is a supermartingale on $\{\mathcal{F}_{n}\}$ and therefore so is the stopped process $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n\wedge\tau}\mathbf{v}\}$ on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$ due to the optional stopping theorem (e.g. Theorem 10.15 of Klenke [2013]). This implies that $\{\mathbf{Y}_{n\wedge\tau}\}$ is a matrix supermartingale on both $\{\mathcal{F}_{n}\}$ and $\{\mathcal{F}_{n\wedge\tau}\}$.
  2. 2.
    For any $\mathbf{v}\in\mathbb{R}^{d}$, we apply the optional sampling theorem (e.g. Theorem 10.11(i,ii) of Klenke [2013]) to the scalar-valued supermartingale $\{\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{n}\mathbf{v}\}$ to obtain that $\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\sigma}\mathbf{v}\geqslant\mathbb{E}(\mathbf{v}^{\mathsf{T}}\mathbf{Y}_{\tau}\mathbf{v}|\mathcal{F}_{\sigma})=\mathbf{v}^{\mathsf{T}}\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})\mathbf{v}$. Therefore $\mathbf{Y}_{\sigma}\succeq\mathbb{E}(\mathbf{Y}_{\tau}|\mathcal{F}_{\sigma})$. ∎
@@
$\text{for any stopping time }\tau,\quad\mathbb{P}(\mathbf{Y}_{\tau}\npreceq\mathbf{A})\leqslant\mathrm{tr}((\mathbb{E}\mathbf{Y}_{0})\mathbf{A}^{-1}).$ (MVI)
@@
We prove the stopped inequality first. Due to Part 3 of Theorem 3.3, $\mathbb{E}\mathbf{Y}_{\tau}\preceq\mathbb{E}\mathbf{Y}_{0}$. Now by (MMI) and the monotonicity of the trace,
@@
As an aside, we present the additional result of a matrix analog of Doob’s $L^{p}$ inequality for forward matrix martingales in Section 9.2. It is not necessarily relevant to our upcoming discussion in statistical contexts but might be of independent interest.